OpenStudy (anonymous):

f(x) = x-x^3 a = -0.2 slope =??

OpenStudy (dumbcow):

could you clarify the question? what does a refer to?

OpenStudy (anonymous):

tangents

OpenStudy (anonymous):

Let f(x) = x-x^3. determine the approximate slope of the line tangent of f at (a, f (a)).

OpenStudy (anonymous):

a = -0.2 slope=?

OpenStudy (anonymous):

I think derivitives

OpenStudy (dumbcow):

ok thank you the slope of f(x) can be determined by taking derivative f'(x) =1 - 3x^2 now plug in x = -0.2 f'(-0.2) = 1 - 3(-.2)^2 = 0.88

OpenStudy (anonymous):

so 0.88 would be the slope?

OpenStudy (dumbcow):

yes f'(x) is same thing as slope at point (x, f(x))

OpenStudy (anonymous):

thank you

OpenStudy (dumbcow):

do you know how to take the derivative of polynomials?

OpenStudy (dumbcow):

d/dx x^n = n*x^(n-1)